OpenStudy (anonymous):

4sin6α/cos^2 3α−sin^2 3α

OpenStudy (anonymous):

Is that \[\frac{4sin6α}{cos^2 3α−sin^2 3α}\]?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

\[\cos^2\theta-\sin^2\theta=\cos(2\theta)\]

OpenStudy (anonymous):

\[\frac{4sin6α}{cos^2 3α−sin^2 3α}=\frac{4sin6α}{cos 6α}=4tan 6a\]